12. Buildz Manufacturing currently produces 2,000 tables per month. The following per unit data for 2,000 tables apply for sales to regular customers: Direct materials $75 Direct manufacturing labor 20 Variable manufacturing overhead 20 Fixed manufacturing overhead 40 Total manufacturing costs $155 What is the per unit cost if 4,000 units are produced assuming the company operates within its relevant range?

A) $125.00 B) $195.00 C) $135.00 D) $155.00 E) None of the above
Answer #1

Variable cost per unit = 75 + 20 + 20 = 115

Fixed cost = 2,000*40 = 80,000

Total Cost = (115*4000) + 80,000 = 540,000

Cost per unit = 540,000/4,000

= 135

Option C is the answer